OpenStudy (anonymous):

Please simplify this: (n+1)^n/(n^n)

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

Well, in order to simplify this, we can first take the whole fraction in a single bracket : \((\cfrac{\sf{n+1}}{\sf{n}})^n\)

mathslover (mathslover):

Is it okay to you?

OpenStudy (anonymous):

What's next?

mathslover (mathslover):

\((\cfrac{n}{n} + \cfrac{1}{n})^n\) .. Can you simplify it further?

OpenStudy (anonymous):

So it's (1+1/n)^n?

mathslover (mathslover):

We can simplify it very easily further too

mathslover (mathslover):

But, in that case , : i) We'll use Binomial Theorem ii) n should be greater than 1

mathslover (mathslover):

You can take it like this : CASE - 1 : If n > 1 \((1+\cfrac{1}{n})^n = 1 + (\cfrac{1}{n} \times n) \\ = 1 +( \cfrac{1}{\cancel{n}^1} \times \cancel{n}^1) \\ = 1+ 1 = 2 \)

mathslover (mathslover):

It is a theorem actually : (1+x)^n = 1 + nx if x is lesser than 1 here, 1/n will be lesser than 1 only if n > 1 so, if n>1 , (1+1/n)^n = 2

OpenStudy (anonymous):

Thanks again.